# MIB: A Mechanistic Interpretability Benchmark

### v1

@@ 5 Related Work
The RAVEL (Huang et al., 2024a) and CausalGym benchmarks (Arora et al., 2024) both enable comparisons across featurization methods, though in more narrow domains. The SAEBench (Karvonen et al., 2025) is similar in concept, though much narrower in scope w.r.t. the kinds of methods that can be evaluated (i.e., only SAEs). Our benchmark compares across a range of tasks, models, and methods.
@@ Acknowledgements
This research was supported by a postdoctoral fellowship under the Zuckerman STEM Leadership Program (A.M.), grants from Open Philanthropy (A.G., Y.B., N.P., D.B.), grants from AI Safety Support Ltd (I.A., R.G.), the OpenAI Superalignment Fast Grant Program (M.H.), the Ariane de Rothschild Women Doctoral Program (D.A.), an armasuisse CYD Doctoral Fellowship (Al.St.), an Azrieli Foundation Early Career Faculty Fellowship (Y.B.), a Google academic gift (Y.B.), and the European Union (ERC, Control-LM, 101165402; Y.B.). The opinions expressed here are those of the authors, and not of any of the mentioned organizations.
@@
Appendix A Limitations
@@
Appendix B Further Details on Materials
@@
B.1 Indirect Object Identification (IOI)
@@ 6 $n_{1}$ $+$ $n_{2}$ = $n_{1}$- $n_{2}$ =
B.3 Multiple-choice question answering (MCQA)
@@
B.4 AI2 Reasoning Challenge (ARC)
By design, each answer in this task is a single token (e.g., A, B, C, D), making the prompt format similar to MCQA (Section B.3).
@@ B.4 AI2 Reasoning Challenge (ARC)
The counterfactual types and generation process are identical to the MCQA counterfactual generation process (described in Section B.3). Due to the varying content of each ARC prompt and the lack of a token-level template between prompts, we include only the format-based counterfactuals, omitting semantic counterfactuals such as “Noun” and “Color” from MCQA. For an example prompt and its counterfactuals, see Table 9.
@@ B.5 Model Performance
Table 12: Model accuracy (0-shot, greedy generation) on ARC counterfactuals.
AP = Answer Position, S = Symbol, RL = Random Letter.
@@ B.6 Leaderboard
For the causal variable localization track, users must provide the hypothesized location of the causal variable $L$, and specify or provide an invertible featurizer function $\mathcal{F}$ via an API format we specify. We will evaluate whether interchange interventions on the features and the variable result in the same behavior for each counterfactual dataset and average the results across them.
Appendix C Details on Circuit Localization Track
C.1 Methods
@@ C.2 Baselines
The loss function comprises two components: (1) a performance metric that measures the discrepancy between the original model’s predictions and the output of the partially ablated model (here, KL divergence); and (2) a regularization term that controls the sparsity of the subgraph determined by the pruning mask. The balance between these components is governed by a hyperparameter $\alpha$. For our experiments, we set $\alpha=10^{-3}$, chosen through a hyperparameter search over $\{10^{-2},10^{-3},\dots,10^{-7}\}$ using a validation set. All other hyperparameters were left at their default values, as specified in Li & Janson (2024).
@@
C.3 Further Circuit Localization Results
@@
Table 13: CPR scores across circuit localization methods and ablation types, and AUROC scores for InterpBench. All evaluations were performed using counterfactual ablations. Higher scores are better. Arithmetic scores are averaged across addition and subtraction; see Table 16 for separate scores. We bold and underline the best and second-best methods per column, respectively.
@@
EActP (CF) 2.28 1.19 - - - 0.92 - - - - -
EAP (mean) 0.31 0.70 0.68 0.98 0.36 0.31 0.14 0.34 0.26 0.34 0.80
EAP (CF) 1.20 0.27 1.28 0.87 0.56 0.75 1.31 0.77 1.08 0.80 0.82
EAP (OA) 0.97 0.71 - - - 0.33 - - - - -
EAP-IG-inputs (CF) 2.01 1.91 3.70 1.90 1.01 1.09 1.55 1.09 1.53 1.04 0.98
EAP-IG-activations (CF) 1.82 1.58 2.09 1.51 0.97 0.78 1.34 1.34 1.70 0.71 0.63
NAP (CF) 0.28 0.30 0.30 0.26 0.27 0.45 1.14 1.36 1.01 0.26 0.26
NAP-IG (CF) 0.78 0.29 1.54 0.43 0.39 0.53 2.03 1.96 1.53 0.26 0.26
IFR 0.72 0.38 0.32 0.16 0.96 0.60 0.35 0.60 0.34 0.36 0.24
UGS 0.96 0.98 - - - 1.18 - - - - -
Table 14: CMD scores for the private test set across circuit localization methods and ablation types (lower is better), and AUROC scores for InterpBench (higher is better). All evaluations were performed using counterfactual ablations. Arithmetic scores are averaged across addition and subtraction. We bold and underline the best and second-best methods per column, respectively.
@@
EActP (CF) 0.01 0.49 - - - 0.55 - - - - -
EAP (mean) 0.26 0.19 0.25 0.04 0.06 0.19 0.18 0.19 0.22 0.19 0.20
EAP (CF) 0.03 0.15 0.06 0.01 0.02 0.08 0.05 0.19 0.04 0.20 0.19
EAP (OA) 0.30 0.16 - - - 0.16 - - - - -
EAP-IG-inp. (CF) 0.02 0.02 0.04 0.00 0.00 0.13 0.03 0.14 0.04 0.10 0.22
EAP-IG-act. (CF) 0.02 0.01 0.03 0.01 0.01 0.05 0.04 0.12 0.04 0.30 0.38
NAP (CF) 0.33 0.31 0.38 0.28 0.26 0.28 0.32 0.32 0.33 0.69 0.69
NAP-IG (CF) 0.22 0.17 0.28 0.17 0.15 0.20 0.24 0.29 0.27 0.67 0.67
IFR 0.30 0.24 0.68 0.82 0.18 0.40 0.65 0.38 0.66 0.63 0.53
UGS 0.04 0.02 - - - - 0.20 - - - -
Table 15: CPR scores for the private test set across circuit localization methods and ablation types, and AUROC scores for InterpBench. All evaluations were performed using counterfactual ablations. Higher scores are better. Arithmetic scores are averaged across addition and subtraction. We bold and underline the best and second-best methods per column, respectively.
@@
EActP (CF) 2.39 1.20 - - - 0.91 - - - - -
EAP (mean) 0.30 0.74 0.65 0.94 0.35 0.32 0.12 0.34 0.26 0.34 0.31
EAP (CF) 1.26 0.28 1.28 0.86 0.56 0.76 1.27 0.78 1.08 0.80 0.26
EAP (OA) 1.11 0.76 - - - 0.34 - - - - -
EAP-IG-inputs (CF) 2.07 1.99 3.40 1.84 1.01 1.08 1.53 1.09 1.53 1.05 0.31
EAP-IG-activations (CF) 1.85 1.61 2.35 1.30 0.97 0.77 1.31 1.40 1.71 0.27 0.28
NAP (CF) 0.27 0.30 0.29 0.26 0.27 0.45 1.10 1.37 1.02 0.26 0.26
NAP-IG (CF) 0.71 0.29 1.44 0.43 0.38 0.53 2.00 2.00 1.54 0.26 0.26
IFR 0.70 0.36 0.32 0.18 0.97 0.60 0.35 0.62 0.34 0.37 0.47
UGS 0.97 1.00 - - - 1.19 - - - - -
@@
Table 16: CPR and CMD scores for Llama-3.1 on arithmetic, separated by operator. Scores are generally similar across operators, and methods follow similar rankings regardless of which operators are used. A notable exception is EAP-IG-activations, where CPR scores are significantly different.
@@
EAP (mean) 0.39 0.06 0.31 0.06
EAP (CF) 0.50 0.02 0.61 0.01
EAP-IG-inputs 0.97 0.00 1.04 0.00
EAP-IG-activations 0.97 0.00 0.97 0.00
NAP (CF) 0.25 0.29 0.29 0.28
NAP-IG (CF) 0.47 0.17 0.40 0.18
IFR 0.91 0.26 0.88 0.19
@@
Appendix D Details on InterpBench Model Training
Faithfulness is a fuzzy and unbound metric. Ideally, we would like to know which edges or nodes are in the circuit in advance, such that we can compute more precise metrics such as precision and recall. Inspired by InterpBench (Gupta et al., 2024), we train a transformer model closely following their methods. The model we use was explicitly trained to predict the indirect objects in the IOI dataset (App. B.1), and implements a simplified version of the IOI circuit described by Gupta et al. (2024).
@@ Appendix D Details on InterpBench Model Training
Appendix E Details on Causal Variable Localization Track
E.1 Causal Abstraction Analysis
@@
E.2 Aligning Unsupervised Features to Causal Variables
@@
E.3 Distributed Alignment Search
@@ E.4 Hyperparameters
E.5 High-level Causal Models and Experimental Details for Each Task
E.5.1 Multiple-choice Question Answering
@@
We conduct two interchange intervention experiments—one for each variable in the causal model—to align the LM’s internal representations with those of the causal model. To align the $X_{\text{Order}}$ variable, we create counterfactual examples where the position of the correct option is altered, while the option label remains the same alphabetically. Conversely, to align the $O_{\text{Answer}}$ variable, we generate counterfactuals in which the position of the correct option is fixed, but the option label is replaced with a different letter.
Figures 8 and 9 show the alignment results of the $X_{\text{Order}}$ variable in the Gemma model using full vector patching and the DAS method, respectively. Both results demonstrate that the $X_{\text{Order}}$ information shifts from the correct symbol token position to the last token position in the middle layers. This behavior aligns with our hypothesized causal model, in which the model first identifies the position of the correct option, before dereferencing it to fetch token value information.
Figure 8: $X_{\text{Order}}$ variable alignment results with full residual vector patching.
Figure 9: $X_{\text{Order}}$ variable alignment results using the subspace identified using the DAS method.
Figures 10 and 11 illustrate the alignment results of the $O_{\text{Answer}}$ variable in the Gemma model, using the full vector patching and DAS methods, respectively. Both results indicate that the $O_{\text{Answer}}$ variable aligns in the later layers—i.e., after the $X_{\text{Order}}$ variable has been established—which is consistent with our hypothesized high-level causal model, where the $O_{\text{Answer}}$ variable derives its information from the $X_{\text{Order}}$ variable.
Figure 10: $O_{\text{Answer}}$ variable alignment results with full vector patching.
Figure 11: $O_{\text{Answer}}$ variable alignment results using the subspace identified using the DAS method.
E.5.2 Arithmetic
We evaluate the hypothesized causal model by conducting an interchange intervention experiment to align the $X_{\text{Carry}}$ variable. The counterfactual examples are designed such that they introduce a carry when the original does not, and remove it when the original includes one. Figures 12 and 13 present the alignment results in the Llama model using full vector patching and the DAS method, respectively. Together, these results suggest that the $X_{\text{Carry}}$ information emerges at the last token position during the middle layers of the model.
Figure 12:
Figure 13:
@@ E.5.2 Arithmetic
E.5.3 RAVEL
[…]
@@ Results.
E.5.4 Indirect Object Indentification

### v2

@@ 5 Related Work
The RAVEL (Huang et al., 2024a) and CausalGym benchmarks (Arora et al., 2024) both enable comparisons across featurization methods, though in more narrow domains. The SAEBench (Karvonen et al., 2025) is similar in concept, though much narrower in scope w.r.t. the kinds of methods that can be evaluated, i.e., only SAEs. Our benchmark compares across a range of tasks, models, and methods.
@@ Acknowledgements
This research was supported by a postdoctoral fellowship under the Zuckerman STEM Leadership Program (A.M.), grants from Open Philanthropy (A.G., Y.B., N.P., D.B.), grants from AI Safety Support Ltd (I.A., R.G.), the OpenAI Superalignment Fast Grant Program (M.H.), the Ariane de Rothschild Women Doctoral Program (D.A.), an armasuisse CYD Doctoral Fellowship (Al.St.), an Azrieli International Postdoctoral Fellowship (A.R.), an Azrieli Foundation Early Career Faculty Fellowship (Y.B.), a Google academic gift (Y.B.), the European Union (ERC, Control-LM, 101165402; Y.B.), and the Apple AIML PhD fellowship (H.O.). The opinions expressed here are those of the authors, and not of any of the mentioned organizations.
@@
Appendix A Table of Notation
Table 4: The notation used throughout the paper, grouped by track.
Track Symbol Meaning
Shared $\mathcal{N}$ The full computation graph; a neural network
Circuit Localization $\mathcal{C}$ A circuit $\in\mathcal{N}$
$u$ A node in the computation graph. Could be an MLP or an attention head
$(u,v)$ An edge from node $u$ to node $v$
$N_{u}$ The number of neurons in node $u$, usually equal to $d_{\text{model}}$
$N_{\mathcal{C}}$ The set of all neurons in $\mathcal{C}$
Causal Variable Localization $\mathcal{H}$ A high-level causal model
$X$ A variable in $\mathcal{H}$
$b$ An base input to the neural network
$\mathcal{D}$ A dataset containing $(b,c)$ pairs
Appendix B Limitations
@@
Appendix C Further Details on Materials
@@
RAVEL 100000 16000 1000
C.1 Indirect Object Identification (IOI)
@@ 6 $n_{1}$ $+$ $n_{2}$ = $n_{1}$- $n_{2}$ =
C.3 Multiple-choice question answering (MCQA)
@@
C.4 AI2 Reasoning Challenge (ARC)
By design, each answer in this task is a single token (e.g., A, B, C, D), making the prompt format similar to MCQA (Section C.3).
@@ C.4 AI2 Reasoning Challenge (ARC)
The counterfactual types and generation process are identical to the MCQA counterfactual generation process (described in Section C.3). Due to the varying content of each ARC prompt and the lack of a token-level template between prompts, we include only the format-based counterfactuals, omitting semantic counterfactuals such as “Noun” and “Color” from MCQA. For an example prompt and its counterfactuals, see Table 10.
@@ C.5 Model Performance
Table 13: Model accuracy (0-shot, greedy generation) on ARC counterfactuals. AP = Answer Position, S = Symbol, RL = Random Letter.
@@ C.6 Leaderboard
For the causal variable localization track, a user must provide an invertible featurizer function $\mathcal{F}$ and token position functions specifying where in an input to apply them. These must be provided as Python scripts. The trained featurizer, inverse featurizer, and token indices must also be provided for each task/model combination. We will evaluate whether interchange interventions on the features and the variable result in the same behavior for each counterfactual dataset and average the results across them.
Appendix D Details on Circuit Localization Track
D.1 Methods
@@ D.2 Baselines
The loss function comprises two components: (1) a performance metric that measures the discrepancy between the original model’s predictions and the output of the partially ablated model (here, KL divergence); and (2) a regularization term that controls the sparsity of the subgraph determined by the pruning mask. The balance between these components is governed by a hyperparameter $\lambda$. For our experiments, we set $\lambda=10^{-3}$, chosen through a hyperparameter search over $\{10^{-2},10^{-3},\dots,10^{-7}\}$ using a validation set. All other hyperparameters were left at their default values, as specified in Li & Janson (2024).
@@
D.3 Further Circuit Localization Results
@@
Table 14: CPR scores across circuit localization methods and ablation types. All evaluations were performed using counterfactual ablations. Higher scores are better. Arithmetic scores are averaged across addition and subtraction; see Table 17 for separate scores. We bold and underline the best and second-best methods per column, respectively.
@@
EActP (CF) 2.30 1.21 - - - 0.85 - - - - -
EAP (mean) 0.29 0.71 0.68 0.98 0.35 0.29 0.33 0.13 0.26 0.34 0.80
EAP (CF) 1.20 0.26 1.29 0.85 0.55 0.85 1.49 1.00 1.08 0.80 0.82
EAP (OA) 0.95 0.70 - - - 0.29 - - - - -
EAP-IG-inputs (CF) 1.85 1.63 3.20 2.08 0.99 1.16 1.64 1.05 1.53 1.04 0.98
EAP-IG-activations (CF) 1.82 1.63 2.07 1.60 0.98 0.77 1.57 0.79 1.70 0.71 0.63
NAP (CF) 0.28 0.30 0.30 0.26 0.27 0.38 1.47 1.69 1.01 0.26 0.26
NAP-IG (CF) 0.76 0.29 1.52 0.42 0.39 0.77 1.71 1.87 1.53 0.26 0.26
IFR 0.58 0.31 0.25 0.09 0.89 0.40 0.38 0.52 0.34 0.36 0.24
UGS 0.97 0.98 - - - 1.17 - - - - -
Table 15: CMD scores for the private test set across circuit localization methods and ablation types (lower is better). All evaluations were performed using counterfactual ablations. Arithmetic scores are averaged across addition and subtraction. We bold and underline the best and second-best methods per column, respectively.
@@
EActP (CF) 0.01 0.48 - - - 0.35 - - - - -
EAP (mean) 0.27 0.24 0.28 0.04 0.07 0.21 0.19 0.17 0.22 0.19 0.20
EAP (CF) 0.03 0.15 0.06 0.01 0.01 0.12 0.08 0.12 0.04 0.20 0.19
EAP (OA) 0.31 0.17 - - - 0.11 - - - - -
EAP-IG-inp. (CF) 0.03 0.02 0.04 0.01 0.01 0.08 0.06 0.14 0.04 0.10 0.22
EAP-IG-act. (CF) 0.02 0.01 0.03 0.01 0.00 0.05 0.07 0.12 0.04 0.30 0.38
NAP (CF) 0.36 0.33 0.40 0.30 0.28 0.24 0.29 0.36 0.33 0.69 0.69
NAP-IG (CF) 0.25 0.19 0.29 0.18 0.17 0.18 0.29 0.33 0.27 0.67 0.67
IFR 0.43 0.70 0.75 0.89 0.22 0.60 0.62 0.49 0.66 0.63 0.53
UGS 0.04 0.02 - - - - 0.19 - - - -
Table 16: CPR scores for the private test set across circuit localization methods and ablation types. All evaluations were performed using counterfactual ablations. Higher scores are better. Arithmetic scores are averaged across addition and subtraction. We bold and underline the best and second-best methods per column, respectively.
@@
EActP (CF) 2.39 1.20 - - - 0.87 - - - - -
EAP (mean) 0.28 0.34 0.64 0.94 0.34 0.29 0.34 0.13 0.26 0.34 0.31
EAP (CF) 1.26 0.27 1.31 0.87 0.54 0.84 1.48 1.06 1.08 0.80 0.26
EAP (OA) 1.07 0.75 - - - 0.29 - - - - -
EAP-IG-inputs (CF) 1.89 1.73 3.03 2.04 0.98 1.61 1.05 0.95 1.53 1.05 0.31
EAP-IG-activations (CF) 1.84 1.61 2.34 1.33 0.98 0.76 1.53 0.83 1.71 0.27 0.28
NAP (CF) 0.27 0.30 0.29 0.25 0.26 0.38 1.46 1.69 1.02 0.26 0.26
NAP-IG (CF) 0.69 0.29 1.42 0.42 0.38 0.77 1.68 1.84 1.54 0.26 0.26
IFR 0.57 0.30 0.25 0.11 0.89 0.40 0.38 0.51 0.34 0.37 0.47
UGS 0.97 1.00 - - - 1.17 - - - - -
@@
Table 17: CPR and CMD scores for Llama-3.1 on the arithmetic public test sets, separated by operator. Scores are generally similar across operators, and methods follow similar rankings regardless of which operators are used. A notable exception is EAP-IG-activations, where CPR scores are significantly different.
@@
EAP (mean) 0.40 0.07 0.31 0.07
EAP (CF) 0.49 0.01 0.61 0.01
EAP-IG-inputs 0.96 0.00 1.03 0.00
EAP-IG-activations 0.98 0.00 0.99 0.00
NAP (CF) 0.26 0.29 0.27 0.27
NAP-IG (CF) 0.43 0.18 0.34 0.18
IFR 0.90 0.24 0.87 0.20
@@
Appendix E Details on InterpBench Model Training
Faithfulness is a fuzzy and unbound metric. Ideally, we would like to know which edges or nodes are in the circuit in advance, such that we can compute more precise metrics such as precision and recall. Inspired by InterpBench (Gupta et al., 2024), we train a transformer model closely following their methods. The model we use was explicitly trained to predict the indirect objects in the IOI dataset (App. C.1), and implements a simplified version of the IOI circuit described by Gupta et al. (2024).
@@ Appendix E Details on InterpBench Model Training
Appendix F Details on Causal Variable Localization Track
F.1 Causal Abstraction Analysis
@@
F.2 Aligning Unsupervised Features to Causal Variables
@@
F.3 Distributed Alignment Search
@@ F.4 Hyperparameters
F.5 High-level Causal Models and Experimental Details for Each Task
F.5.1 Multiple-choice Question Answering
@@
We highlight two interchange intervention experiments—one for each variable in the causal model—to align the LM’s internal representations with those of the causal model. For both variables, we use counterfactuals in which the position of the correct option and the option label is replaced with a different letter. (To be clear, in the main text, we report the average across this counterfactual and counterfactuals that only change position and only change the option label.)
Figures 8(a) and 8(b) show the alignment results of the $X_{\text{Order}}$ variable in the Gemma model using full vector patching and the DAS method, respectively. This result shows that we can use DAS isolate the answer position in the middle layers above the correct label token and then the final token. This behavior aligns with our hypothesized causal model, in which the model first identifies the position of the correct option, before dereferencing it to fetch token value information.
Figures 8(c) and 8(d) illustrate the alignment results of the $O_{\text{Answer}}$ variable in the Gemma model, using the full vector patching and DAS methods, respectively. The success of full vector patching at the last layer of the last token is logically guaranteed, as this is where the model predicts the next token. However, full vector patching fails at other locations because the output token is tangled with ordering information and the positional encoding. On the other hand, DAS is able to disentangle the output token from this other information at basically every layer.
(a) $X_{\text{Order}}$ variable alignment results with full residual vector patching.
(b) $X_{\text{Order}}$ variable alignment results using the subspace identified using the DAS method.
(c) $O_{\text{Answer}}$ variable alignment results with full vector patching.
(d) $O_{\text{Answer}}$ variable alignment results using the subspace identified using the DAS method.
@@
(a) $X_{\text{Carry}}$ variable alignment results in Llama with full vector patching and counterfactuals that differ only by the carry the one variable.
(b) $X_{\text{Carry}}$ variable alignment results in Llama using the DAS method and counterfactuals that differ only by the carry the one variable.
(c) $X_{\text{Carry}}$ variable alignment results in Llama with full vector patching and counterfactuals that are randomly sampled.
(d) $X_{\text{Carry}}$ variable alignment results in Llama with the DAS method and counterfactuals that are randomly sampled.
F.5.2 Arithmetic
We evaluate the hypothesized causal model by conducting an interchange intervention experiment to align the $X_{\text{Carry}}$ variable. The counterfactual examples are designed such that they introduce a carry when the original does not, and remove it when the original includes one. Figures 10(a) and 10(b) present the alignment results in the Llama model using full vector patching and the DAS method, respectively. At first, these results suggest that the $X_{\text{Carry}}$ information emerges at the last token position during the middle layers of the model. However, this is likely a representation of the tens-digit output, which perfectly correlates with the ones-carry variable in these counterfactual. Indeed, when we use random counterfactuals in which the output is entirely independent from the carry the one value, we see a total failure.
@@ F.5.3 RAVEL
The Llama model achieves 66.5% accuracy on the task and the Gemma model achieves 70.5% accuracy on the task. The dataset contains many cities, so considering that there are many countries and languages to choose from, this is good performance. We filter out failure cases.
@@
Results.
[…]
LogitDifferenceSubjectTokenSubjectPositionTextPrompt×0.768×2.005+0.048
(a) General causal model for indirect object identification
LogitDifference = 3.16SubjectToken = 1SubjectPosition = 1 TextPrompt: ”As Carl and Maria left the consulate, Carl gave a fridge to” ×0.768×2.005+0.295
(b) Specific example showing prediction for ”Carl”
Figure 14: Causal model for indirect object identification. The TextPrompt is processed to extract the subject token $S_{\text{Tok}}$ and the subject position $S_{\text{Pos}}$. The output variable mechanism (1) compares the token and position to the input, and determines whether the token and position were inverted or the same and sets PositionSignal and TokenSignal to 1 and -1 accordingly and (2) computes $0.048+0.768\texttt{TokenSignal}+2.005\texttt{PositionSignal}$, with position having a stronger influence than token identity.
F.5.4 Indirect Object Indentification